Reduce the PDE uxx+xuyy
into a canonical form in the region
x<0.
1
Expert's answer
2021-07-16T08:00:14-0400

Comparing the given DE with

"aU_{xx} + 2bU_{xy} + cU_{yy}"

We have that

"a = 1, b=0, c = x"

But

"\dfrac{dy}{dx} = \dfrac{b \pm \sqrt{b^2-ac}}{a}"

In this case,

"\dfrac{dy}{dx} = \pm \sqrt{-x}"

This value will be a real number since we are considering the region x < 0

"dy = \pm (-x)^\frac{1}{2}dx \\ y = \pm \dfrac{ 2(-x)^\frac{3}{2}}{3} + c"

Set

"\xi = 3y -2 (-x)^\frac{3}{2} ....... (*)"

And

"\eta = 3y + 2 (-x)^\frac{3}{2}...... (**)"

"\xi_x = -3(-x)^{\frac{1}{2}} \\
\eta_x = 3(-x)^{\frac{1}{2}}"

"\xi_y = 3 \\ \eta_y = 3"

"\xi_{xx} =- \frac{3}{2}(-x)^{-\frac{1}{2}} \\
\eta_{xx}= \frac{3}{2}(-x)^{-\frac{1}{2}}"

"\xi_{yy}= \eta_{yy} = 0"

So we have that

"U_x = U_{\xi} \xi_x + U_{\eta} \eta_{x} = -3(-x)^{\frac{1}{2}} U_{\xi} + 3(-x)^{\frac{1}{2}} U_{\eta}"

"U_{xx} = -\frac{3}{2}(-x)^{-\frac{1}{2}} U_{\xi}-9x U _{\xi \xi} + \frac{3}{2}(-x)^{-\frac{1}{2}} U_{\eta}-9x U _{\eta \eta}"

"U_y = U_{\xi} \xi_y + U_{\eta} \eta_{y} = 3 U_{\xi} + 3 U_{\eta}"

"U_{yy} = 9U_{\xi \xi} + 3 U_{\eta \eta}"

From (*) and (**), we have that

"\eta - \xi = 4 (-x)^{\frac{3}{2}} \\ (\dfrac{4}{\eta - \xi})^ \frac{1}{3} = (-x)^{-\frac{1}{2}} .... (***)"

Substituting the value of "U_{xx}" and "U_{yy}" into the equation, we have

"U_{xx} + xU_{yy} = (U_{\eta} - U_{\xi}) \frac{3}{2}(-x)^{-\frac{1}{2}}"

And substituting (***), we have

"U_{xx}+xU_{yy} = \dfrac{3(U_{\eta} - U_{\xi})}{[2(\eta - \xi)]^ \frac{1}{3}}"